ANNUAL REPORT OF THE CENTRAL BUREAU FOR ASTRONOMICAL TELEGRAMS

1988

A total of 175 Circulars were issued during 1988, down from the record 230 of the preceding year, but significantly more than enough to put 1988 in second place. A large number of the Circulars reported continuing observations of SN 1987A in the LMC, but many were also due to residual activity from the record 16 discoveries and 17 recoveries of comets announced in 1987.

Nevertheless, with 26 new discoveries of supernovae 1988 was the first year the alphabet was exhausted in real time—although 1954 still holds the record when objects recognized much later are included—and while there were only three recoveries of periodic comets, the number of new comets given provisional designations, 15, was second only to 1987. Considering also two objects observed in 1987 but not announced until 1988, there were no fewer than seven detections of new sungrazing comets with the coronagraph/polarimeter on the orbiting Solar Maximum Mission. Two other comets, 1988e and 1988g, represent the first definite example of ordinary long-period comets traveling in the same orbit.

Among the other items mentioned on the Circulars during the year were two optical and one x-ray galactic nova and two novae in the Large Magellanic Cloud. An optical counterpart was discovered to the eclipsing millisecond pulsar PSR 1957+20, and binary millisecond pulsars were reported in 47 Tucanae. There were for the first time confirmed observations of an occultation by Pluto and the detection of an anomalous brightening of (2060) Chiron.

As noted in the 1987 report, the availability of SPAN dramatically increased the impact of the Bureau’s Computer Service, allowing remote logins for reading the Circulars as soon as they are issued. In July 1988 arrangements were made to mail the Circulars electronically to Computer Service subscribers who could conveniently be reached by SPAN or BITNET. These subscribers are warned that e-mail frequently tends to disappear permanently or temporarily into a black hole and that it is impossible for the Bureau to undertake to provide electronic replacements other than by having the subscribers log in to the Computer Service to obtain the issues they are missing.

As was also noted last year, the increased dissemination of the computerized Circulars, both in terms of the number of subscribers and the frequency of issue, is continuing to reduce the need for the Bureau’s telegram service. The total of 35 ‘telegram books’ was therefore quite small—but in fact identical with the total in 1985. In September arrangements were completed with Western Union for the receipt of incoming telexes and telegrams on one of the MicroVAXes used by the Bureau. This has the advantage that all messages are thereby available in machine-readable form, but the elimination of the old mechanical TWX machines means that it is now not always quite so obvious that a message has arrived. Outgoing telegrams and telexes are also now handled in a more straightforward manner. The telegram service continues to be essential for astronomers in the second world and most of the third world, although there has been some progress in that the Computer Service has been accessed from India, and e-mail distribution of the Circulars is possible to Argentina, Brazil and Chile.

Since the beginning of April the copy for the printed Circulars has been prepared using TEX and a laser printer, replacing the computer and mechanical printer introduced just 5\frac{1}{2} years earlier. The widespread emphasis nowadays on electronic distribution tends to offset this improvement, but it should be noted that the printed Circulars remain the “official” version.

Daniel W. E. Green attended to much of the scientific operation of the Bureau during the year and in October was promoted from Assistant to Associate Director. Assistance with orbit computations was also provided by Assistant Director Conrad M. Bardwell and by Syuichi Nakano. Peter Davies and his colleagues at IPSO continued to help save costs by redisseminating telegrams to Australia and New Zealand. Donna Thompson again attended to the Circulars accounts and has provided general secretarial assistance (although Green attends to the telegram accounts), and additional clerical assistance was provided by Paul K. Dzus. The number of subscribers to the printed Circulars peaked at 754 in April and diminished to only 711 at the end of the year. On the other hand, by the end of the year the number of subscribers to the Computer Service had increased to almost 150, of whom 60 percent were receiving the Circulars directly by e-mail.

Brian G. Marsden Director of the Bureau